OpenStudy (anonymous):

How many grams of iron (III) chloride must decompose to produce 78.4 milliliters of iron metal, if the density of iron is 7.87 g/mL? Show all steps of your calculation as well as the final answer. FeCl3 → Fe + Cl2

OpenStudy (aaronq):

you need to balance the equation first, because you need to use the stoichiometric coefficients.

OpenStudy (aaronq):

once you do that you need to find how many moles of Fe correspond to the given quantity: "78.4 milliliters of iron metal, if the density of iron is 7.87 g/mL" first find the mass: \[density(\rho)=\frac{ mass(m) }{Volume(V) }\] (equation 1) then find the moles (n): \[moles(n)= \frac{ mass(m) }{ Molar mass(M) }\] (equation 2) Once you find the moles of Fe, find how much FeCl3 you need. Since everything in stoichiometry works in moles, you need to find the moles of Fe equivalent to FeCl3 using the stoichiometric coefficients. If the equation is this: aFeCl3 → bFe + Cl2 (a,b,c are the stoichiometric coefficients) you can build a relationship using their ratios: \[\frac{ n _{FeCl _{3}} }{ a } =\frac{ n_{Fe}}{ b }\] after you found the moles of FeCl3, convert to mass using equation 2.

OpenStudy (anonymous):

Thank You So Much!!